OpenStudy (anonymous):

Solve e-2(x+1) = 2. Round your answer to the nearest hundredth. Answer x ≈ -1.35 x ≈ -1.45 x ≈ 1.65 x ≈ 1.78

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

e-2x-2=2 2x=2.73-4 x=-0.635

OpenStudy (anonymous):

So C?

OpenStudy (ghazi):

so B